OpenStudy (anonymous):

Find all solutions 7 sin^2x-14 sin x+2=-5 Please help!!

OpenStudy (campbell_st):

so start be equating the equation to zero and then you basically have a quadratic \[7\sin^2(x) - 14\sin(x) + 7 = 0\] factor out the 7 \[7(\sin^2(x) - 2\sin(x) + 1) = 0\] which now looks like a perfect square inside the brackets hope it helps

OpenStudy (anonymous):

Would that be my final answer or can i keep going?

OpenStudy (campbell_st):

well when you factor you get \[7(\sin(x) -1)^2 = 0\] so then you need to find the zero what makes sin(x) - 1 = 0